Your university is running a special offer on tuition. This year's tuition cost is $16,000. Next year's tuition cost is scheduled to be $16,640. The university offers to discount next year's tuition at a rate of 4% if you agree to pay both years' tuition in full today. How much is the total tuition bill today if you take the offer?
Answer #1

PRESENT VALUE = FUTURE VALUE/(1+r)

TUITION COST IS PAYABLE IN THE BEGINNING OF THE YEAR. SO

16000 IS PAYABLE TODAY AND 16640 IS PAYABLE AFTER A YEAR. IF WE PAY TODAY BOTH YEARS FEES :

TOTAL TUITION BILL TODAY = 16000 + (16640/(1+0.04)) = 16000 + 16000 = 32000

ANSWER : $32000 (Thumbs up please)